OpenStudy (anonymous):

what is the sum of the square roots of 16i?

OpenStudy (anonymous):

\[i \sqrt{16}= \pm 4i\]

OpenStudy (anonymous):

i think @casanovaxx has done it incorrectly

OpenStudy (anonymous):

answer should be (4*\[e^{ipi/4}\])

OpenStudy (anonymous):

and off course both +/- sign

OpenStudy (anonymous):

Okay, the possible answers given are -8, 0, 4\[\sqrt{2}\] +4\[\sqrt{2i}\], and 8i

OpenStudy (anonymous):

sorry that third one is , \[4\sqrt{2}+4\sqrt{2i}\]

OpenStudy (anonymous):

8i? or the correction i added for the long one?

OpenStudy (anonymous):

square roots of \(i\) are \(\frac{\sqrt{2}}{2}+\frac{\sqrt{2}}{2}i\) and \(-\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}i\) and get 0

OpenStudy (anonymous):

the 16 of course is unimportant, you multiply the above numbers by 4 when you add you still get zero

OpenStudy (anonymous):